# Matthias Schonlau

### v1

@@ 1 Introduction
Leonardo da Vinci wrote, a poet would struggle to “describe with words what a painter is able to [depict] in an instant” (Janson and Janson, 2004). But not everybody can paint like da Vinci. Likewise, it is often easier to write a few pages of text than to construct a sensible graph. Arguably, readers appreciate graphs more than text. Empirical evidence suggests, graphs even lend an aura of scientific credibility (Tal, 2015). For those of us who cannot paint like da Vinci, software makes it easier to create graphs. In this paper we introduce software for creating hammock plots.
@@
Figure 1: A scatter plot and a hammock plot of repair record (1-5 “stars”) vs car origin from the auto data. With only two axes, the usefulness of the hammock plot may not yet be apparent.
a scatter plot of the variables repair record and car origin (foreign vs domestic) and a corresponding hammock plot. The repair record takes the values 1 through 5 corresponding to 1 through 5 “stars”. To reduce overplotting, the scatter plot uses jittering, i.e., we add a little bit of random noise. A scatter plot is not the most effective way to display these two variables; our interest is in the layout of the axes. In a scatter plot the two axes are aligned at a right angle. A hammock plot uses parallel coordinates: the axis for repair record is parallel to the axis for car origin. On the axes, there are bars with univariate statistics; the height of the bar is proportional to the number of observations it contains. Axes are connected by bars. The width of a connecting bar is also proportional to the number of observations a bar contains.
@@
Figure 2: Hammock plot of the auto data. Foreign cars are highlighted in orange. A category for missing values, where present, is displayed near the bottom. Here, the dashed black line was added to show the separation between the areas for missing and non-missing values.
The graph consists of columns with univariate information and lines/ bars that connect the columns. We discuss each in turn.
Each variable is shown in a column or axis; the variable name is displayed beneath the column. For categorical variables, each column contains a bar that corresponds to one category. The length of the bar is proportional to the number of observations it represents. For example, the variable foreign in the last column has two categories, “foreign” and “domestic” representing foreign and domestic cars. We see there are more than twice as many domestic cars as foreign cars in this data set. For numerical variables, there is also a “bar” for each value. However, for values occurring only once, the bar is so thin that it looks like a line. So, the univariate display for a numerical value looks like a rug plot (see the discussion in Cox (2025)), which we typically see at the bottom of a density plots to show the individual data points of a continuous variable.
@@
Neighboring axes are connected by bars. In Figure 2, the connecting bars are rectangles, but other shapes could be used (e.g., parallelograms, which are faster to compute). The width of the rectangles is proportional to the number of observations it contains. Rectangles containing very few observations appear as a line. The rectangles each represent a cell of the two-way contingency table of the two adjacent variables, i.e., they represent bivariate statistics.
Figure 2 highlights foreign cars in orange. We can follow the observations belonging to foreign cars throughout the graph. For example, we see that foreign cars tend to have less headroom, a better repair record, lower weight, and higher mileage. For the continuous variables mileage (miles per gallon) and weight we can also make a statement about correlations. Mileage and weight are negatively correlated because the lines cross in opposite directions: high values of mileage correspond to low values of weight and conversely. In fact, the correlation between mileage and weight turns out to be 0.80.8-0.8- 0.8.
The remainder of this paper is as follows. Section 2 gives a brief history of the hammock plot and similar plots that followed. Section 3 gives numerous examples and explains the Stata syntax in the context of the examples. Section 4 concludes.
@@ 2 A brief history of the hammock plot and related plots
Perhaps the best known plot among the hammock-type plots is the alluvial plot (Rosvall and Bergstrom, 2010). Whereas the hammock plot uses bars (parallelograms or rectangles), alluvial plots use rounded curves for the bivariate elements that connect axes. These plots were initially only used to visualize network variables over time. Alluvial plots are now also used to visualize categorical variables; they do not accommodate numerical variables. In Stata, alluvial plots are implemented in the package alluvial Naqvi [2024a].
A similar looking plot, the Sankey diagram (Sankey, 1898; Schmidt, 2008), visualizes flows of energy and materials rather than variables. A historical Sankey diagram is shown in Figure 3.
Figure 3: A Sankey diagram of flows of energy (Sankey, 1898)
A Sankey diagram may have circular flows or one flow may stop in the middle of the graph. The Sankey diagram is implemented in Stata (Naqvi, 2024b). The names “Sankey diagram” and “Alluvial plot” are sometimes conflated in the literature. If a plot visualizes variables, it is not a Sankey diagram. A more detailed overview over all of these plots can be found in Schonlau (2024). A Stata implementation of the parallel coordinate plot is in parplot (Cox, 2007).
3 Examples and Syntax
We will introduce the syntax by example. Our first example features a mixture of numerical and categorical variables, highlighting, and missing values (Section 3.1). In the following section, Section 3.2, we trace individual observations in a dataset related to the Tour de France. The variables in this example all use the same scale (seconds). The samescale option shows the effect of standardizing the layout across axes. Section 3.3, contains an example where our interest lies in connecting bars between axes; the univariate bars are largely redundant and are omitted. Section 3.4 shows the other extreme: parallel univariate bars without connectors. The examples include a graphical representation of socio-demographic table, enriching or replacing the typical “Table 1” in empirical studies.
@@ 3.1 Highlighting
The hammock plot in Figure 2 highlighted foreign cars. It was created with the following command:
@@
    . label var mpg mileage
    . label var weight weight
    . label var headroom headroom
@@
    > yline(10)
The option missing adds room at the bottom across all variables for a category for any missing values. We highlight the variable and specific values of that variable with hivar and hival. The label on the plot is not the number highlighted (1), but its label (“foreign”). The option label adds the labels and the univariate bars. Likewise, the variable name shown below each axis is the variable label. So, we see the label “mileage” rather than the variable name “mpg”. We can highlight (hival(0)) a single value, missing values (hival(.)), a range (hival(>27)) or several values in different colors (e.g., hival(1 4 5)). It is also possible to specify an external variable for highlighting, i.e., a variable that is not among the variables in the list of variables shown in the graph.
@@
Figure 2 relied on the default colors. We can specify different colors in the option colorlist. Figure 4) shows the auto data where cars with a large gas mileage are highlighted (specifically, mileage>27). We see cars with high mileage are cheaper, they tend to have lower weight, a better repair record (rep78), and less headroom.
@@
    > colorlist(gs11 magenta )
The first color in option colorlist is the default color; any subsequence colors are used for highlighting in the order specified in hival(). You can list all named colors in Stata with graph query colorstyle; it is also possible to specify colors in RGB format.
We can also include colors created from the palette command. The palette command saves colors in RGB format (e.g., ”0 0 255”) in r(p) . We can input this directly into the colorlist option
@@ 3.2 Tracing individual observations
The main group of riders is called the peloton. In bicycling, it is advantageous to ride in groups to draft off other riders. You can see the peloton across the stages. Additional groups may form as riders try to break out ahead of the peloton (break-away groups and chase-groups) or cannot keep up (the so-called “autobus” group). In particular, in the mountain stages several groups may form. Mountain stages are those with variable names ending in an “m”.
Stage 20 is a time trial indicated by the variable name ending with a “t”. In a time trial, riders ride individually and cannot draft off their teammates. There are no groups in the time trial.
@@
Figure 5: Hammock plot of the first and the last stages of the tour de France data. The eventual winner, Pogacar, is highlighted. This plot scales each axis individually from fastest to slowest rider. Times are not comparable across axes.
@@
We removed the labels (nolabel) to avoid clutter. barwidth(.2) reduces the width of the bars connecting the axes (relative to the default barwidth(1)). Sometimes connecting bars of individual observations are barely visible because the bar is too thin. minbarfreq(5) increases the width of any connecting bars with fewer than 5 observations as if they had 5 observations. That is, bars with 5 or fewer observations all have the same width. Also note the variable used for highlighting, rider, is not among the variables plotted.
Figure 5 nicely shows fast and slow riders. However, a unit of time is shown differently on each parallel axis: the difference between the slowest time (at the top) and the fastest time (at the bottom, above the missing values) is not constant across stages. We subtract the time of the fastest rider for each stage and denote the new variables d1-d21. These new variables represent how many seconds have passed since the stage winner crossed the finishing line. We can then use the option samescale(varlist) to enforce the same scale for all variables. The revised plot is shown in Figure 6.
Figure 6: Hammock plot of time lost to stage winners for multiple stages. The higher a rider on the plot, the more time he lost. All axes are on the same scale. The eventual winner, Pogacar, is highlighted.
@@
For the revised plot, Figure 6 , we first computed the d1-d21 variables by subtracting the fastest stage time. The variable name is changed from s‘i’ to ‘i’ with d‘i’ ranging from 1 to 21. THe code is as follows:
@@ 3.3 A connectors-only Hammock plot
Sometimes it makes sense to remove the univariate charts and concentrate on the connectors. Figure 7 shows marital status over 5 time points (or waves) in the GESIS panel (https://www.gesis.org/en/gesis-panel), a survey panel in Germany. A survey panel is a longitudinal survey where the same individuals (or units) are surveyed repeatedly over time. The most common category in Figure 7 is “married” (living together rather than apart), followed by “single”. We can see that marital status does not change much within a few short years. What might surprise is the small subset of observations that went from “married” to “single” (rather than “divorced” or one of the other categories). It turns out that marital status is self-assessed and some respondent preferred “single” over “divorced” or another category.
@@
For comparison, Figure 8 shows a plot that includes the univariate bars. Figure 7 was created as follows:
@@
Figure 8: Comparison plot to Figure 7 with univariate bars included.
nounibar requests not to plot bars on the axes. space(0.1) leaves some space between the connectors. We specified nolabel because the same label applies for all variables. Instead, we added the labels by hand using Stata’s text option which is passed through to the plotting command. The x-coordinate for the text is always $11$ because the text is for the first x-variable and the first x-variable has the x-coordinate $11$. The y-coordinates are spaced between 0 and 100; the exact values are determined by trial and error. (Temporary optional arguments such as yline(10 20 30 40 50) can help finding desired coordinates). barwidth(3) triples the width of the connecting bars. We add some space on the left and right using Stata’s graphregion(margin()) option. This avoids that some text runs off the page. For variety, we specified a different color using a color intensity of 30%. We did not reserve an area for missing values in this plot.
More on the size and shape of bars
uni_fraction(0.5) controls the length of the univariate bars. By default, 50% of the vertical space is filled with univariate bars. If any two bars overlap, this fraction can be adjusted downward to avoid the overlap. This fraction can also be adjusted upward to suit. Likewise, barwidth(1) can be adjusted to control the width of the connectors to reduce visual clutter or to reveal more detail. The number specified is an adjustment factor relative to the default, 1 (see an example in Figure 7). Connecting bars that represent a very small fraction of observations can be so thin that they are barely visible or do not render properly. As seen in the code for Figure 5, option minbarfreq(1) specifies that a rectangle must represent at least minbarfreq observations. All bars corresponding to fewer than minbarfreq observations are displayed as if they contained minbarfreq observations. The default value, 1, does nothing because all rectangles contain at least 1 observation.

### v2

@@ 1 Introduction
Leonardo da Vinci wrote, a poet would struggle to “describe with words what a painter is able to [depict] in an instant” (Janson and Janson, 2004). But not everybody can paint like da Vinci. Likewise, it is often easier to write a few pages of text than to construct a sensible graph. Arguably, readers appreciate graphs more than text. Empirical evidence suggests graphs even lend an aura of scientific credibility (Tal, 2015). For those of us who cannot paint like da Vinci, software makes it easier to create graphs. In this paper we introduce a Stata command for creating hammock plots.
@@
Figure 1: A scatter plot and a corresponding hammock plot of repair record (1-5 “stars”) vs car origin from the auto data. The hammock plot has parallel axes. Univariate bars (on the axes) are proportional to frequency of the corresponding category. Connectors (between the axes) are proportional to the corresponding bivariate frequencies. The frequency of missing connectors (e.g. repair record “1” to “foreign”) is zero. With only two axes, the usefulness of the hammock plot may not yet be apparent.
a scatter plot of the variables repair record and car origin (foreign vs domestic) and a corresponding hammock plot. The repair record takes the values 1 through 5 corresponding to 1 through 5 “stars”. To reduce overplotting, the scatter plot uses jittering, i.e., we add a little bit of random noise. In a scatter plot the two axes are aligned at a right angle. A hammock plot uses parallel coordinates: the axis for repair record is parallel to the axis for car origin. On the axes, there are univariate bars. They display univariate statistics: the height of the bar is proportional to the number of observations it contains. Axes are connected by connectors. The width of a connector is also proportional to the number of observations it contains.
The connectors represent a visualization of the two-way contingency table of repair record and car origin:
  • . tab rep78 foreign, cell nofreq
        repair          Car origin
        record      Domestic    Foreign         Total
             1          2.90       0.00          2.90
             2         11.59       0.00         11.59
             3         39.13       4.35         43.48
             4         13.04      13.04         26.09
             5          2.90      13.04         15.94
         Total         69.57      30.43        100.00
There are eight connectors between the two axes and eight non-zero cells in the two-way contingency table. The bars on the axes visualize the marginal statistics of the two variables shown in the “Total” row and column of the contingency table.
@@
Figure 2: Hammock plot of the auto data. Foreign cars are highlighted. A category for missing values, where present, is displayed near the bottom. Here, the black line was added to show the separation between the areas for missing and non-missing values.
The graph consists of axes with univariate information (univariate bars) and connectors that connect the axes. We discuss each in turn.
Each variable is shown on a parallel axis; the variable name is displayed beneath the axes. For categorical variables, each axis contains a bar that corresponds to one category. The height of the bar is proportional to the number of observations it represents. For example, the variable foreign in the last axis has two categories, “foreign” and “domestic” representing foreign and domestic cars. We see there are more than twice as many domestic cars as foreign cars in this data set. For numerical variables, there is also a “bar” for each value. However, for values occurring only once, the bar is so thin that it looks like a line. So, the univariate display for a numerical value looks like a rug plot (see the discussion in Cox (2025)), which we typically see at the bottom of a density plots to show the individual data points of a continuous variable. Each axis is scaled individually from minimum to maximum of the corresponding variable. For example, the axis for mileage ranges from 12 to 41; that of weight from 1760 to 4840.
@@
Neighboring axes are connected by connectors. In Figure 2, the connectors are rectangles, but other shapes could be used (e.g., parallelograms, which are faster to compute). The width of the connector is proportional to the number of observations it contains. Connectors containing very few observations appear as a line (even though they are just thin connectors). The connectors each represent a cell of the two-way contingency table of the two adjacent variables, i.e., they represent bivariate statistics.
Figure 2 highlights foreign cars in a different color. We can follow the observations belonging to foreign cars throughout the graph. For example, we see that foreign cars tend to have less headroom, a better repair record, lower weight, and higher mileage. For the continuous variables mileage (miles per gallon) and weight we can also make a statement about correlations. Mileage and weight are negatively correlated because the connectors cross in opposite directions: high values of mileage correspond to low values of weight and conversely. In fact, the correlation between mileage and weight turns out to be 0.8-0.8.
The remainder of this paper is as follows. Section 2 gives a brief history of the hammock plot and similar plots that followed. Section 3 gives an overview over syntax and options. Section 4 gives numerous examples and explains the Stata syntax in the context of the examples. Section 5 concludes.
@@ 2 A brief history of the hammock plot and related plots
Perhaps the best known plot among the hammock-type plots is the alluvial plot (Rosvall and Bergstrom, 2010). Whereas the hammock plot uses parallelograms or rectangles to connect axes, alluvial plots use rounded curves. These plots were initially only used to visualize network variables over time. Alluvial plots are now also used to visualize categorical variables; they do not accommodate numerical variables. In Stata, alluvial plots are implemented in the package alluvial Naqvi [2024a].
A similar looking plot, the Sankey diagram (Sankey, 1898; Schmidt, 2008), visualizes flows of energy and materials rather than variables. A Sankey diagram may have circular flows or one flow may stop in the middle of the graph. The Sankey diagram is implemented in Stata (Naqvi, 2024b). The names “Sankey diagram” and “Alluvial plot” are sometimes conflated in the literature. If a plot visualizes variables, it is not a Sankey diagram. A more detailed overview over all of these plots can be found in Schonlau (2024). A Stata implementation of the parallel coordinate plot is in parplot (Cox, 19 March 2007).
3 Syntax
hammock varlist [if] [in] , [options]
The following sections list options alongside a brief description. As in the helpfile, we underline abbreviations for options. For a more detailed description please consult the helpfile. For the most recent version of this command see the Boston SSC archive (ssc install hammock) or my github page (Schonlau, 2025).
@@ 3.1 Highlighting
hivariable(varname)
Name of the variable to highlight.
hivalues(string)
List of values of hivariable to highlight.
colorlist(str)
Default color and colors used for highlighting.
uni_colorlist(str)
Default color and colors used for highlighting univariate bars.
3.2 Layout of univariate bars
nounibar
Do not show univariate bars.
uni_fraction(real)
Proportion of vertical space allocated to univariate bars.
3.3 Layout of labels
nolabel
Do not show value labels.
label_methodlist(str)
Specify individual labeling methods for each variable.
label_too_many(real)
When a variable has at least label_too_many unique values, show only the minimum and maximum label.
label_min_dist(real)
For method min_dist, show only labels that are at least label_min_dist apart.
labelopt(str)
Options passed to added_text_options, for example to control label text size.
label_format(str)
Display format of numeric labels.
3.4 Layout of labels
nolabel
Do not show value labels.
label_min_dist(real)
Specify the minimum distance between two labels on the same axis.
labelopt(str)
Options passed to added_text_options, for example to control label text size.
label_format(str)
Display format for numerical labels.
3.5 Missing values
missing
Display missing values.
missing_fraction(real)
Proportion of vertical space allocated to missing values.
3.6 Layout of connectors between axes
barwidth(real)
Change the width of connectors to reduce clutter.
minbarfreq(int)
Specify the minimum connector width.
shape(str)
Shape of the connectors: parallelogram (or par) or rectangle (default).
outline
Outline the edges of semi-translucent connectors (rarely needed).
3.7 Other options
space(real)
Control the fraction of space allocated to labels and univariate bars relative to connectors.
samescale(varlist)
Use the same axis scale for each variable specified.
subspace(real)
Adjust empty space between univariate bars and connectors (rarely needed).
aspectratio(real)
Aspect ratio of the plot region (rarely needed).
graph_options
Additional options passed through to graph, twoway.
4 Examples
Our first example features a mixture of numerical and categorical variables, highlighting, and missing values (Section 4.1). In the following section, Section 4.2, we trace individual observations in a dataset related to the Tour de France. The variables in this example all use the same scale (seconds). The samescale option shows the effect of standardizing the layout across axes. Section 4.3, contains an example where our interest lies in connectors between axes; the univariate bars are largely redundant and are omitted. Section 4.4 shows the other extreme: parallel univariate bars without connectors. The examples include a graphical representation of socio-demographic table, enriching or replacing the typical “Table 1” in empirical studies.
4.1 Highlighting
The hammock plot in Figure 2 highlighted foreign cars. The print version of this figure (and all figures) is in grayscale; the online version is in color. The color version was created with the following command:
@@ 4.1 Highlighting
    . label var mpg "mileage"
    . label var weight "weight"
    . label var headroom "headroom"
@@
    >  yline(8) ///
    >  colorlist(ltblue sandb)
The grayscale version replaces the colors using colorlist(gs12 gs8). The hammock plot displays variable labels (when present) and the display looks best with short labels. We therefore label some variables.
The option missing adds room at the bottom across all variables for a category for any missing values. We highlight the variable and specific values of that variable with hivar and hival. The label on the plot is not the number highlighted (1), but its label (“foreign”). Likewise, the variable name shown below each axis is the variable label. So, we see the label “mileage” rather than the variable name “mpg”. We can highlight (hival(0)) a single value, missing values (hival(.)), a range (hival(>27)) or several values in different colors (e.g., hival(1 4 5)). It is also possible to specify an external variable for highlighting, i.e., a variable that is not among the variables in the list of variables shown in the graph.
@@
Rather than relying on default colors, we can specify colors in the option colorlist. Figure 3) shows the auto data where cars with a high gas mileage are highlighted (specifically, mileage>27). We see cars with high mileage are cheaper, they tend to have lower weight, a better repair record (rep78), and less headroom.
@@
    > colorlist(gs11 sandb)
The first color in option colorlist is the default color; any subsequence colors are used for highlighting in the order specified in hival(). In the grayscale version of this Figure, the colorlist is replaced with colorlist(gs12 gs8). You can list all named colors in Stata with graph query colorstyle; it is also possible to specify colors in RGB format.
We can also include colors created from the user-written colorpalette command (Jann, 2023). The colorpalette command saves colors in RGB format (e.g., ”0 0 255”) in r(p) . We can input this directly into the colorlist option
@@
Specifying Labels
Figure 3) used default labeling: add labels for all values unless there are too many (>8>8). In that case, show only labels for the minimum and maximum.
We can change the default behavior either by changing what number of values is “too many” in option label_\_too_\_many, or by specifying individual labels for each axis. For the 6 variables in Figure 3, we could specify the labeling method for each variable using label_\_methodlist(none minmax all all all minmax). This would remove the labels of the first variable, display labels for minimum and maximum for the second and sixth variable, and displaying all labels for the remaining 3 variables.
@@ 4.2 Tracing individual observations
The main group of riders is called the peloton. In bicycling, it is advantageous to ride in groups to draft off other riders. You can see the peloton across the stages. Additional groups may form as riders try to break out ahead of the peloton (break-away groups and chase-groups) or cannot keep up (the so-called “autobus” group). In particular, in the mountain stages several groups may form. Mountain stages are those with variable labels ending in an “m”.
Stage 20 is a time trial indicated by the variable label ending with a “t”. In a time trial, riders ride individually and cannot draft off their teammates. There are no groups in the time trial.
@@
Figure 4: Hammock plot of the first and the last stages of the tour de France data. The eventual winner, Pogacar, is highlighted. Each axis is scaled individually from fastest to slowest rider. Times are not comparable across axes.
@@
We removed the labels on the axes (nolabel) to avoid clutter. barwidth(.2) reduces the width of the connectors (relative to the default barwidth(1)). Sometimes connectors of individual observations are barely visible because the connector is too thin. minbarfreq(5) increases the width of any connectors with fewer than 5 observations as if they had 5 observations. That is, connectors with 5 or fewer observations all have the same width. Also note the variable used for highlighting, rider, is not among the variables plotted. uni_fraction(.15) scales the bars on the axes to cover 15% of the available vertical space. The remaining 85% is white. This option can be adjusted to avoid overlap between the bars. Some text is added (“slow”, “fast”, “missing”). The text is placed with y-values between 0-100 (here: 97, 10, 3) and with the x-value 1 corresponding to the first axis.
Figure 4 nicely shows fast and slow riders. However, a unit of time is shown differently on each parallel axis: the difference between the slowest time (at the top) and the fastest time (at the bottom, above the missing values) is not constant across stages; some stages are much longer than others. We subtract the time of the fastest rider for each stage and denote the new variables d1-d21. These new variables represent how many seconds have passed since the stage winner crossed the finishing line. We can then use the option samescale(varlist) to enforce the same scale for all variables. The revised plot is shown in Figure 5.
Figure 5: Hammock plot of time lost to stage winners for multiple stages. The higher a rider on the plot, the more time he lost. All axes are on the same scale. The eventufual winner, Pogacar, is highlighted.
@@
For the revised plot, Figure 5 , we first computed the d1-d21 variables by subtracting the fastest stage time. The new variables are named d‘i’ (rather than s‘i’) with ‘i’ ranging from 1 to 21. The variables are created first:
@@ 4.3 A connectors-only Hammock plot
A hammock plot consists of alternating displays of univariate statistics (bars) and bivariate statistics (connectors). When concentrating on the transitions, it sometimes makes sense to only show the connectors.
Figure 6: Marital status over 5 waves of a survey panel. This hammock plot is identical to Figure 7 except that here univariate bars are included.
Figure 6 shows marital status over 5 time points (or waves) in the GESIS panel (https://www.gesis.org/en/gesis-panel), a survey panel in Germany. A panel survey is a longitudinal survey where the same individuals (or units) are surveyed repeatedly over time. Here, both the bars and the connectors are shown. The most common category in Figure 7 is “married” (living together rather than apart), followed by “single”. We can see that marital status does not change much within a few short years. The connectors look very similar to the univariate bars; the univariate bars take up space without adding information.
Figure 7 shows the same plot but removes the univariate bars.
@@
Now we can more easily concentrate on the transitions from wave to wave. What might surprise is the small subset of observations that went from “married” to “single” (rather than “divorced” or one of the other categories). It turns out that marital status is self-assessed and some respondent preferred “single” over “divorced” or another category.
Figure 7 was created as follows:
@@
As before, the grayscale version of this Figure can be created by replacing
colorlist(blue%30) with colorlist(gs8). nounibar requests not to plot univariate bars on the axes. space(0.1) leaves some space between the connectors. We specified nolabel because the same label applies for all variables. Instead, we added the labels by hand using Stata’s text option which is passed through to the plotting command. minbar(20) requests that each connector has a minimum thickness (the minimum thickness is specified in terms of the number of respondents it represents, here 20 respondents). The x-coordinate for the text is always $1$ because the text is for the first x-variable and the first x-variable has the x-coordinate $1$. The y-coordinates are spaced between 0 and 100; the exact values are determined by trial and error. (Temporary optional arguments such as yline(10 20 30 40 50) can help finding desired coordinates). barwidth(3) triples the width of the connectors. We add some space on the left and right using Stata’s graphregion(margin()) option. This avoids that some text runs off the page. For variety, we specified a different color using a color intensity of 30%. We did not reserve an area for missing values in this plot.
More on the size and shape of bars and connectors
uni_fraction(0.5) controls the height of the univariate bars. By default, 50% of the vertical space is filled with univariate bars. If any two bars overlap, this fraction can be adjusted downward to avoid the overlap. This fraction can also be adjusted upward to suit. Likewise, barwidth(1) can be adjusted to control the width of the connectors to reduce visual clutter or to reveal more detail. The number specified is an adjustment factor relative to the default, 1 (see an example in Figure 7). connectors that represent a very small fraction of observations can be so thin that they are barely visible or do not render properly. As seen in the code for Figure 4, option minbarfreq(1) specifies that a connector must represent at least minbarfreq observations. All bars corresponding to fewer than minbarfreq observations are displayed as if they contained minbarfreq observations. The default value, 1, does nothing because all connectors contain at least 1 observation.